Jefferson Lab HALL A Logbook for October 07 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 211043 10/31/07 23:51 Chiranjib TARGET----->>>> H2
  • 211038 10/31/07 23:39 Chiranjib TARGET------->>> Fe
  • 211031 10/31/07 23:24 Chiranjib TARGET------>>> C12
  • 211028 10/31/07 23:18 cisbani Pb Momentum acceptance looks bad
  • 211025 10/31/07 23:11 Chiranjib TARGET ------>>>> DUMMY
  • 211024 10/31/07 23:03 cisbani beam down for a while
  • 211021 10/31/07 23:01 Chiranjib Loop 3/ Loop 1
  • 211018 10/31/07 22:52 Chiranjib TARGET------>>> H2
  • 211013 10/31/07 22:39 Chiranjib TARGET ---->>> Pb
  • 211008 10/31/07 22:26 cisbani new right momentum = 290.738 MeV/c
  • 211007 10/31/07 22:25 cisbani new left momentum 291.710 MeV/c
  • 211006 10/31/07 22:22 Chiranjib TARGET ---->>> He
  • 211003 10/31/07 22:06 Chiranjib TARGET ------>>>>Fe
  • 211000 10/31/07 21:55 Chiranjib TARGET ----->>> C12
  • 210997 10/31/07 21:42 Chiranjib TARGET------>>> DUMMY
  • 210994 10/31/07 21:31 Chiranjib TARGET SCREENS II
  • 210993 10/31/07 21:30 Chiranjib TARGET SCREENS I
  • 210992 10/31/07 21:26 Chiranjib TARGET------>>> H2
  • 210988 10/31/07 21:17 brads BPM/BCM stripchart for longer period
  • 210986 10/31/07 21:13 Chiranjib TARGET------->>>Pb
  • 210981 10/31/07 20:57 Chiranjib TARGET ----->>>> He
  • 210980 10/31/07 20:51 cisbani Correction: left arm mometum at 235.648
  • 210979 10/31/07 20:46 cisbani LEft arm mometum at 252.741
  • 210976 10/31/07 20:43 cisbani Carbon target: sort of halo in raster
  • 210975 10/31/07 20:42 Chiranjib TARGET ---->> Fe
  • 210970 10/31/07 20:32 cisbani BPM / BCM stripcharts
  • 210969 10/31/07 20:31 cisbani ratiation stripchart
  • 210967 10/31/07 20:30 Adaq Compton and OTR Stripchart
  • 210965 10/31/07 20:28 Chiranjib TARGET ------>>> C12
  • 210964 10/31/07 20:25 cisbani Hall A Tools
  • 210961 10/31/07 20:21 cisbani beam back
  • 210958 10/31/07 20:10 Chiranjib TARGET ALARM RESOLVED
  • 210957 10/31/07 19:23 Target Alarm ALARM
  • 210956 10/31/07 19:18 cisbani Right Momentum at 334.448 MeV/c
  • 210955 10/31/07 18:48 Chiranjib TARGET----->>> DUMMY
  • 210952 10/31/07 18:32 Chiranjib TARGET------>>> H2
  • 210947 10/31/07 18:21 cisbani Left arm momentum at 252.741 MeV/c
  • 210946 10/31/07 18:19 Chiranjib TARGET ---->>> Pb
  • 210943 10/31/07 18:04 Chiranjib TARGET ----->>>> He
  • 210938 10/31/07 17:45 Chiranjib TARGET------>>>> Fe
  • 210933 10/31/07 17:34 Chiranjib TARGET------>>> C12
  • 210924 10/31/07 17:16 Chiranjib TARGET ------->>>> DUMM
  • 210919 10/31/07 17:03 Chiranjib Target ----->>>> Pb
  • 210914 10/31/07 16:46 Chiranjib Target------->>> He ( Loop 1)
  • 210913 10/31/07 16:42 cisbani Left Arm Mometum to 271.075
  • 210908 10/31/07 16:02 ron gilman shift summary
  • 210902 10/31/07 15:51 hyao target moved from Loop3-Pb15cm to Loop1-H-15cm
  • 210901 10/31/07 15:50 hyao target moved from Loop3-Pb15cm to Loop1-He-10cm
  • 210888 10/31/07 15:13 hyao target moved from dummy-10cm to Loop3-Pb15cm
  • 210885 10/31/07 15:01 hyao target moved from Beam-right-C to dummy-10cm
  • 210884 10/31/07 14:58 brads Modified runstart.tcl target list to match target ladder labels
  • 210876 10/31/07 14:46 gilman cut sfor right arm online physics replay
  • 210874 10/31/07 14:42 hyao target moved from Beam-right-Fe to Beam-right-C
  • 210869 10/31/07 14:25 fansler Right Q1 Current is Railing
  • 210868 10/31/07 14:19 hyao target moved from Loop3-Pb15cm to Beam-right-Fe
  • 210867 10/31/07 14:12 hyao target moved from dummy-10cm to Loop3-Pb15cm
  • 210862 10/31/07 14:01 hyao target moved from emety to dummy-10cm
  • 210861 10/31/07 13:44 LeRose Right Dipole
  • 210857 10/31/07 13:12 gilman Hall A Tools screen
  • 210856 10/31/07 12:13 V. Sulkosky New LHRS HVs for S1 and S2
  • 210855 10/31/07 12:12 brads Tech: Right dipole NMR won't enable
  • 210854 10/31/07 12:01 V. Sulkosky S1 and S2 ADCs After HV Changes
  • 210853 10/31/07 11:55 gilman requested beam
  • 210848 10/31/07 11:15 A. Camsonne NaI work
  • 210844 10/31/07 08:57 hyao target moved from Beam Right Iron to Empty
  • 210843 10/31/07 08:30 brads Rebooted iocha48 (Right Q1 ioc)
  • 210842 10/31/07 08:26 gilman shur down for tour
  • 210838 10/31/07 08:05 Z. Meziani Shift Summary
  • 210830 10/31/07 07:33 brads OTR behavior changed around 0400
  • 210823 10/31/07 07:11 meziani Snapshot of hall A tools
  • 210820 10/31/07 06:55 meziani spot_L
  • 210816 10/31/07 06:32 meziani Pb at 358.708 MeV/c repeat of run # 1536
  • 210784 10/31/07 03:24 jpchen raster size during yesterday morning
  • 210782 10/31/07 03:17 jpchen y-targ for Pb and Fe at P0=474.672
  • 210779 10/31/07 03:06 jpchen y_targ for Pb and Fe at P0=412.6
  • 210766 10/31/07 02:12 Z. Meziani Spot_Left results
  • 210757 10/31/07 01:43 brads Tech: Troubleshooting activity on the Right magnets
  • 210756 10/31/07 01:40 Z. Meziani Spot_L
  • 210751 10/31/07 01:27 Z.Meziani Snapshot results for Harp scan B
  • 210750 10/31/07 01:26 Z.meziani Snapshot for harpscan results for A
  • 210749 10/31/07 01:08 meziani Harp scan results snapshot
  • 210748 10/31/07 01:05 brads Right Q3 hall probe working
  • 210747 10/31/07 01:02 brads Tech: RHRS Q1 LHe reading back 'inf' on Tools GUI
  • 210745 10/31/07 00:13 E.Chudakov Swing shift summary
  • Tuesday

  • 210744 10/30/07 23:58 Shabestari Target charts
  • 210743 10/30/07 23:21 brads Right dipole power supply in strange state
  • 210742 10/30/07 22:06 E.Chudakov Control access
  • 210741 10/30/07 21:57 brads Right Q3 hall probe history -- flatlined Monday 2pm
  • 210740 10/30/07 21:47 brads Right Q3 hall probe is somehow disabled?
  • 210731 10/30/07 21:26 E.Chudakov Tools at 442 MeV
  • 210730 10/30/07 21:25 Shabestari Target motions @ P0 = 474.672 MeV/c
  • 210729 10/30/07 21:25 brads Restarted Compton DAQ
  • 210726 10/30/07 21:18 E.Chudakov Right Q3
  • 210719 10/30/07 20:58 E.Chudakov RHRS Q3 at 442 MeV
  • 210702 10/30/07 19:42 V. Sulkosky Update to onlineGUI software
  • 210699 10/30/07 19:35 V. Sulkosky Added Raster and BPM Plots to Detector Replay.
  • 210696 10/30/07 19:26 jpchen Pb and LH2 rates
  • 210691 10/30/07 19:16 jpchen Fe data do not show a clear peak either
  • 210688 10/30/07 19:11 E.Chudakov Tools, BCM screens
  • 210685 10/30/07 18:58 E.Chudakov Raster size
  • 210679 10/30/07 18:34 jpchen No peak in y_targ for Pb target run this morning!
  • 210674 10/30/07 16:00 gilman day shift summary
  • 210653 10/30/07 13:59 gilman spot check of beam
  • 210640 10/30/07 13:43 hyao Target moved from empty to dummy-10cm
  • 210639 10/30/07 13:36 gilman harp scan after beam returns
  • 210638 10/30/07 13:24 R. Michaels Scaler clock mystery resolved
  • 210637 10/30/07 13:21 A. Camsonne NaI check
  • 210636 10/30/07 12:44 Meekins Target UPS prblems
  • 210635 10/30/07 11:58 brads Opportunistic controlled access during last half of beam studies
  • 210634 10/30/07 08:45 hyao Target moved from Loop3-15cm+Pb to empty
  • 210631 10/30/07 08:08 Eric Voutier Shift summary
  • 210627 10/30/07 07:39 brads Restarted Compton DAQ
  • 210626 10/30/07 07:30 meziani Snapshots of loop1 & 3
  • 210625 10/30/07 07:28 meziani Target sudden power glitch
  • 210604 10/30/07 04:13 meziani Target charts snapshot
  • 210603 10/30/07 04:09 Eric Voutier Beam strip chart
  • 210602 10/30/07 03:57 Eric Voutier Beam problems
  • 210597 10/30/07 03:05 A. Camsonne NaI status
  • 210594 10/30/07 03:03 A. Camsonne NaI status
  • 210593 10/30/07 03:02 A. Camsonne NaI status
  • 210587 10/30/07 02:31 jpchen T1/T3 scalers difference still there
  • 210585 10/30/07 02:25 jpchen raster check with spot_L
  • 210584 10/30/07 02:22 jpchen target vertical scan
  • 210559 10/30/07 01:13 jpchen Lift encoder position updated
  • 210552 10/30/07 01:03 brads Good information, but please don't use external links for data
  • 210551 10/30/07 00:55 Eric Voutier OTR check
  • 210550 10/30/07 00:54 A. Camsonne NaI Status
  • 210549 10/30/07 00:53 A. Camsonne NaI Status
  • 210543 10/30/07 00:11 A. Camsonne Added NaI HV to start of run
  • Monday

  • 210541 10/29/07 23:59 E.Chudakov Swing shift summary
  • 210536 10/29/07 23:44 brads Make sure energy and position FFB is enabled before starting tests
  • 210531 10/29/07 23:12 A. Camsonne NaI HV changed
  • 210530 10/29/07 22:25 E.Chudakov The magnets are OK
  • 210529 10/29/07 22:03 E.Chudakov Cryo glitch
  • 210524 10/29/07 21:51 A. Camsonne NaI HV changed
  • 210523 10/29/07 21:50 R. Michaels I could be wrong about clock
  • 210522 10/29/07 21:38 V. Sulkosky BPM Pedestals
  • 210521 10/29/07 21:35 spiegel TECHS
  • 210520 10/29/07 21:32 R. Michaels clock discrepancy
  • 210519 10/29/07 21:16 R. Michaels bugfix in ~a-onl last_scaler
  • 210516 10/29/07 20:38 R. Michaels various trigger issues
  • 210515 10/29/07 20:28 atrig Trigger-Download
  • 210514 10/29/07 20:09 atrig Trigger-Download
  • 210513 10/29/07 20:08 atrig Trigger-Download
  • 210512 10/29/07 20:07 atrig Trigger-Download
  • 210511 10/29/07 19:58 V. Sulkosky Cryo Alarm in Hall A
  • 210510 10/29/07 19:49 V. Sulkosky Updated Angle Info in Hall A Tools
  • 210509 10/29/07 19:09 V. Sulkosky Online Replay Instructions
  • 210508 10/29/07 19:08 V. Sulkosky Online Replay Instructions
  • 210498 10/29/07 18:08 LeRose Right Dipole "Quench"
  • 210496 10/29/07 17:21 A. Camsonne NaI work
  • 210491 10/29/07 17:08 ycwang Target Screens
  • 210490 10/29/07 16:54 ycwang Target screens
  • 210489 10/29/07 16:33 V. Sulkosky S2 signal switched at Discriminator
  • 210488 10/29/07 16:31 gilman day shift summary
  • 210486 10/29/07 16:20 V. Sulkosky S2 Right Arm ADC R-1 and R-3
  • 210485 10/29/07 16:15 V. Sulkosky LHRS and RHRS BPMA readout should be fixed
  • 210484 10/29/07 15:38 brads/sulkos LHRS BPMA readout should be fixed
  • 210457 10/29/07 13:39 R. Michaels R-arm EDTM repair
  • 210456 10/29/07 13:38 R. Michaels trig. cleanup, redundant scaler reads
  • 210450 10/29/07 13:17 atrig Trigger-Download
  • 210448 10/29/07 13:12 atrig Trigger-Download
  • 210447 10/29/07 12:49 LeRose Right Dipole "Quench"
  • 210446 10/29/07 12:33 brads Floor angle labels refreshed around 90 and 120 degrees
  • 210445 10/29/07 12:29 brads NaI bubbler filled with mineral oil
  • 210417 10/29/07 10:41 A. Camsonne NaI HV change
  • 210413 10/29/07 10:32 A. Camsonne HV increased NaI
  • 210412 10/29/07 10:12 gilman left arm moved to ~89 degrees
  • 210411 10/29/07 09:21 gilman will try to move left arm
  • 210410 10/29/07 09:19 gilman right arm moved out to nearly 90 degrees
  • 210407 10/29/07 08:44 jpchen 10/29 owl shift summary
  • 210402 10/29/07 08:24 gilman calorimeter HV dropped to 600 V
  • 210398 10/29/07 08:07 gilman target changed
  • 210396 10/29/07 07:48 jpchen Left HRS motion problem.
  • 210393 10/29/07 07:31 brads OTR stripchart snap
  • 210391 10/29/07 07:19 jpchen NaI HV changed to 1000 V
  • 210388 10/29/07 06:57 K.Allada Target ROC rebooted
  • 210382 10/29/07 05:26 meziani right arm preshower ADC 40 and shower ADC 55
  • 210381 10/29/07 05:20 jpchen Additional data taking and NaI check before energy change
  • 210376 10/29/07 04:47 jpchen Right Dipole quench
  • 210375 10/29/07 04:44 meziani S2 Right Arm ADC R-1 and R-3
  • 210372 10/29/07 04:37 meziani Bubbler for NaI detector gas flow
  • 210357 10/29/07 03:10 meziani Spot_R
  • 210356 10/29/07 03:08 meziani spot_L
  • 210354 10/29/07 02:56 jpchen P changed to 285.87 MeV/c
  • 210351 10/29/07 02:36 Z.-E. Mezian Cherenkov left arm
  • 210350 10/29/07 02:34 jpchen L/R trigger rate difference
  • 210347 10/29/07 02:27 jpchen BPM /BCM stripe charts
  • 210346 10/29/07 02:25 jpchen ozone stripe chart
  • 210345 10/29/07 02:24 jpchen Hall A tools screen
  • 210342 10/29/07 02:20 Z.-E. Mezian S2 R-PMTs ADC 1 and ADC 3
  • 210324 10/29/07 00:05 Ramesh Shift Summary
  • 210321 10/29/07 00:01 R. Michaels Re: trig in scalers.
  • 210320 10/29/07 00:00 Shabestari Target:
  • Sunday

  • 210319 10/28/07 23:59 Shabestari Target screen
  • 210312 10/28/07 23:05 brads Scaler rate stripcharts for Runs 20214, 1390 (He target)
  • 210309 10/28/07 22:56 Yoomin NaI bubbler web cam and fan monitor--howto
  • 210305 10/28/07 22:12 brads OTR x pos a function of beam current(!?)
  • 210295 10/28/07 21:07 Ramesh Hall A Ozone level and Compton strip charts
  • 210294 10/28/07 21:02 Ramesh BPMs and Hall A tools
  • 210281 10/28/07 20:09 R. Michaels re: trigger rates
  • 210273 10/28/07 19:22 brads Strange trigger rates
  • 210261 10/28/07 17:54 brads Please do NOT close the Alarm handler
  • 210254 10/28/07 17:31 brads Restarted Left arm CODA
  • 210253 10/28/07 17:30 brads Run 1408 is junk run
  • 210234 10/28/07 16:06 Solvignon Shift summary
  • 210232 10/28/07 15:58 Shabestari Target screen
  • 210211 10/28/07 14:15 Solvignon Kinematic change: P=0.526 GeV/C
  • 210193 10/28/07 13:15 Solvignon SpotR
  • 210192 10/28/07 13:12 Solvignon spotL
  • 210188 10/28/07 12:51 Solvignon BCM stripchart
  • 210187 10/28/07 12:49 Solvignon BPM stripchart
  • 210186 10/28/07 12:46 Solvignon Compton stripchart
  • 210185 10/28/07 12:46 xqian target status
  • 210184 10/28/07 12:45 Solvignon Ozone stripchart
  • 210183 10/28/07 12:44 Solvignon Hall A tools
  • 210174 10/28/07 11:47 Solvignon Kinematic change: P=0.565 GeV/c
  • 210171 10/28/07 11:38 Solvignon CSR NaI detector checks
  • 210158 10/28/07 11:15 Solvignon Left arm S2 TDC
  • 210155 10/28/07 11:06 Solvignon Left cerenkov
  • 210154 10/28/07 11:04 Solvignon Right cerenkov
  • 210149 10/28/07 10:53 V. Sulkosky BPMA raster data missing/broken in both arms
  • 210144 10/28/07 10:31 brads Compton ROC5 rebooted this morning
  • 210141 10/28/07 10:30 brads Raster correction isn't critical -- we'll fix BPMA raster signal next access
  • 210140 10/28/07 10:23 brads BPMA raster data missing/broken in both arms
  • 210135 10/28/07 09:47 Solvignon Kinematic change: P=0.606 GeV/c
  • 210126 10/28/07 09:13 brads OTR stripchart
  • 210114 10/28/07 08:10 Z.-E. Mezian Shift Summary
  • 210097 10/28/07 07:19 Yoomin spot_L
  • 210096 10/28/07 07:12 Yoomin spot_R
  • 210088 10/28/07 06:48 Meziani Compton/ROC 5 problem
  • 210076 10/28/07 05:41 Meziani Ozone/Radiation charts
  • 210075 10/28/07 05:38 Meziani Target charts
  • 210074 10/28/07 05:33 Z.-E. Mezian Snapshot of Hall A General tools
  • 210053 10/28/07 03:55 Meziani Momentum setting
  • 210044 10/28/07 03:32 jpchen vacuum alarm limit change
  • 210039 10/28/07 03:02 A. Camsonne HV lowered
  • 210024 10/28/07 02:03 jpchen scattering chamber vacuum alarm
  • 210016 10/28/07 01:28 Meziani Fast Feedback
  • 210015 10/28/07 01:27 jpchen scatteirng chamber vacuum slowing improving
  • 210014 10/28/07 01:21 jpchen scattering chamber vacuum 2.7*10-5
  • 210012 10/28/07 01:12 jpchen heat load
  • 210009 10/28/07 00:58 Yoomin target loop1
  • 210004 10/28/07 00:37 Yoomin target moved
  • 210001 10/28/07 00:29 brads OTR seems to be working now
  • 210000 10/28/07 00:14 A. Camsonne NaI HV
  • 209997 10/28/07 00:02 Ramesh Shift Summary
  • Saturday

  • 209995 10/27/07 23:51 Chiranjib TARGET SCREEN with Fe(left)
  • 209992 10/27/07 23:50 Chiranjib TARGET ---->>> Fe (left)
  • 209985 10/27/07 23:01 V. Sulkosky Left BPM/Raster II
  • 209983 10/27/07 22:39 brads Compton rate too high? (6kHz/uA)
  • 209982 10/27/07 22:14 Ramesh OTR in
  • 209978 10/27/07 21:49 V. Sulkosky Left Arm Gas Cherenkov gas turned to "Low Flow"
  • 209977 10/27/07 21:43 V. Sulkosky LHRS BPM/Raster Problem Probably Fixed
  • 209976 10/27/07 21:30 A. Camsonne Gate NaI
  • 209975 10/27/07 21:21 R. Michaels re: L-arm BPM/raster
  • 209972 10/27/07 21:05 A. Camsonne HV lowered
  • 209971 10/27/07 21:04 brads Beam center position is (0, 1.5) +- 0.2 mm
  • 209970 10/27/07 21:04 A. Camsonne Cable 96 recrimped
  • 209969 10/27/07 20:52 Ramesh OTR out (will put back in)
  • 209966 10/27/07 20:08 brads Low fluid level in NaI Bubbler
  • 209964 10/27/07 20:03 Chiranjib TARGET----->>>C12( left)
  • 209961 10/27/07 19:52 Chiranjib CORRECTION
  • 209959 10/27/07 19:51 brads General Tools snapshot
  • 209958 10/27/07 19:48 V. Sulkosky Fastbus Extra Hits LHRS
  • 209957 10/27/07 19:30 Chiranjib PAD RUNS
  • 209954 10/27/07 19:29 brads Opportunistic Controlled Access
  • 209951 10/27/07 18:20 V. Sulkosky BPM/Raster Left arm Problem
  • 209950 10/27/07 17:59 Chiranjib TAREGT SCREENS II
  • 209949 10/27/07 17:58 Chiranjib TARGET SCREENS I
  • 209948 10/27/07 17:45 Chiranjib TARGET ------>>> C12
  • 209947 10/27/07 17:39 Ramesh Both hrs's at 60 deg and at 0.98952 GeV/c
  • 209944 10/27/07 17:26 Chiranjib TARGET------->>> LOOP1
  • 209943 10/27/07 17:22 Ramesh Left hrs at 60 deg
  • 209937 10/27/07 16:14 Seonho Choi About NaI High Voltage
  • 209936 10/27/07 16:06 Seonho Choi Shift Summary (Day 10/27/2007)
  • 209931 10/27/07 15:49 xinhu Yan Target moved to Loop3
  • 209923 10/27/07 15:04 Xinhu Yan Target moved to Loop1
  • 209917 10/27/07 14:04 Seonho Choi LHRS at 30 degrees, beam current raised to 50uA
  • 209911 10/27/07 13:11 R. Michaels skipping /adaql1/data1 for now
  • 209903 10/27/07 12:18 Seonho Choi Pedestal subtraction for LHRS etc.
  • 209900 10/27/07 11:50 Seonho Choi NaI HV adjustment
  • 209895 10/27/07 11:05 Seonho Choi /adaql1/data1 disk degraded mode
  • 209889 10/27/07 10:47 A. Camsonne NaI HV change
  • 209888 10/27/07 10:46 A. Camsonne Pedestal subtraction off on left arm
  • 209870 10/27/07 09:36 A. Camsonne NaI changed HV
  • 209867 10/27/07 09:31 Seonho Choi Preshower/Shower HV change aborted
  • 209866 10/27/07 09:09 Seonho Choi RHRS Preshower/Shower HV
  • 209851 10/27/07 08:11 Xinhu Yan Target Status
  • 209848 10/27/07 08:03 Z.E. Meziani shift summary
  • 209840 10/27/07 06:57 A. Camsonne NaI HV change
  • 209814 10/27/07 04:33 jpchen Reduced Left Gas Cherenkov HVs, reasonable now.
  • 209809 10/27/07 04:12 Yoomin Correction for run R#20160, L#1293
  • 209807 10/27/07 04:06 Yoomin RHRS Run#20160, LHRS Run#1293--Collimator Open
  • 209800 10/27/07 03:35 jpchen Left Gas Cherenkov still not workingg
  • 209790 10/27/07 02:46 R. Michaels restarted auto abu
  • 209789 10/27/07 02:34 jpchen optics study
  • 209766 10/27/07 00:43 Meziani DAQ
  • 209763 10/27/07 00:08 Ramesh Shift Summary
  • 209762 10/27/07 00:07 Ramesh Bull's eye scan result
  • 209761 10/27/07 00:01 Kalyan Target ladder position offset in encoder
  • Friday

  • 209759 10/26/07 23:52 Ramesh BTA not working
  • 209739 10/26/07 23:23 Kalyan target to optics
  • 209738 10/26/07 23:23 Kalyan Results of Target Ladder Scan
  • 209737 10/26/07 23:00 brads Harp scan display from the Accelerator menu (Monticello)
  • 209730 10/26/07 21:39 Ramesh Bull's eye scan point (2,2)
  • 209726 10/26/07 21:21 Ramesh BPM positions at (0,0) for bull's eye scan
  • 209718 10/26/07 20:37 Y. Oh / A. C NaI work
  • 209713 10/26/07 17:25 brads Ion chambers calibrated
  • 209712 10/26/07 17:09 K.Allada target empty
  • 209711 10/26/07 17:02 Xinhu Yan updated NaI pedstals
  • 209706 10/26/07 16:13 averett Day Shift Summary
  • 209704 10/26/07 15:56 LeRose Left arm controls whiteout/rebooted
  • 209703 10/26/07 12:45 LeRose Target Screens
  • 209702 10/26/07 12:17 brads Run Plan for Friday
  • 209701 10/26/07 12:05 brads Compton strip charts not working
  • 209700 10/26/07 11:57 averett HRS-L cerenkov gas flow
  • 209699 10/26/07 11:57 brads RHRS low momentum settling time
  • 209698 10/26/07 11:55 Jack Left Arm Gas Cherenkov gas turned on
  • 209697 10/26/07 11:13 brads Fixed NaI channels 229, 155, 72
  • 209688 10/26/07 10:13 R. Michaels T2 "fix" and L-arm gas C
  • 209687 10/26/07 10:13 V. Sulkosky NaI Channel 155
  • 209686 10/26/07 09:27 atrig Trigger-Download
  • 209685 10/26/07 09:26 atrig Trigger-Download
  • 209684 10/26/07 08:51 V. Sulkosky NaI Channel 255
  • 209683 10/26/07 08:51 V. Sulkosky NaI Channel 255
  • 209682 10/26/07 08:10 R. Michaels no need to repeatedly download trigger (and T2)
  • 209681 10/26/07 08:04 V. Sulkosky Owl Shift Summary
  • 209679 10/26/07 06:32 V. Sulkosky LHRS S1 and S2 ADC spectra
  • 209678 10/26/07 05:57 jpchen target screens
  • 209677 10/26/07 05:52 jpchen HV adjustments for S1/S2
  • 209676 10/26/07 05:42 V. Sulkosky No ADC signal for Cerenkov (LHRS)
  • 209673 10/26/07 05:27 V. Sulkosky Left and Right DAQs Hung
  • 209668 10/26/07 03:51 V. Sulkosky T2 rate zero after Trigger Download
  • 209667 10/26/07 03:34 jpchen beam tuning hitting EMPTY target frame (dummy).
  • 209666 10/26/07 03:27 V. Sulkosky Beam hitting Dummy Target
  • 209665 10/26/07 03:12 atrig Trigger-Download
  • 209664 10/26/07 02:46 atrig Trigger-Download
  • 209659 10/26/07 02:36 atrig Trigger-Download
  • 209650 10/26/07 01:47 jpchen Right spectrometer to 15 degree
  • 209649 10/26/07 01:36 A. Camsonne Compton rate display
  • 209648 10/26/07 01:31 V. Sulkosky Hall A in Beam Permit
  • 209642 10/26/07 01:07 V. Sulkosky Hall in Controlled Access
  • 209636 10/26/07 00:47 brads NaI update for Run 1256
  • Thursday

  • 209632 10/25/07 23:41 Ramesh NaI problematic channels
  • 209629 10/25/07 23:12 brads Brief status update...
  • 209628 10/25/07 22:54 brads Runplan for FRIDAY owl --> Day shift
  • 209627 10/25/07 22:48 brads Run plan for Thursday OWL through DAY shift
  • 209626 10/25/07 22:47 E.Chudakov HRSR shower channel #49 is fixed
  • 209623 10/25/07 21:23 Chiranjib TARGET SCREENS III
  • 209622 10/25/07 21:21 Chiranjib TARGET SCREENS II
  • 209621 10/25/07 21:20 Chiranjib TARGET SCREENS I
  • 209616 10/25/07 20:04 A. Camsonne NaI channel 17 DC offset
  • 209615 10/25/07 19:59 R. Michaels L-arm trigger restored, EDTM not working
  • 209610 10/25/07 19:35 A. Camsonne NaI Channel 252 moved
  • 209605 10/25/07 18:45 Haiyan Gao Shift Summary
  • 209602 10/25/07 18:21 A. Camsonne Pedestal check
  • 209595 10/25/07 17:58 atrig Trigger-Download
  • 209594 10/25/07 17:57 brads Bubbler web-cam
  • 209593 10/25/07 17:48 brads NaI fan monitoring instructions
  • 209587 10/25/07 16:52 brads LHRS HVs back on, alarm handler masks removed
  • 209578 10/25/07 16:41 brads LHRS in the hut (late-entry) + misc update
  • 209575 10/25/07 16:04 Holmstrom Shower ADC 48
  • 209561 10/25/07 15:09 fansler Check list
  • 209560 10/25/07 15:02 atrig Trigger-Download
  • 209559 10/25/07 14:38 jpchen target loop 1refilled to 172 psi.
  • 209558 10/25/07 12:04 Holmstrom Cryo Crate crash
  • 209557 10/25/07 11:57 R. Michaels pin polarity
  • 209556 10/25/07 10:38 V. Sulkosky LHRS BPM/Raster
  • 209555 10/25/07 10:35 V. Sulkosky LHRS BPM/Raster Cable
  • 209554 10/25/07 09:32 brads Striptool config files in /home/hacuser/Setups/StripTool/e05-110/
  • 209553 10/25/07 08:52 brads Re-enabled the tripped HVs in Beamline crate, slot 1
  • 209551 10/25/07 08:01 Eric Voutier Shift Summary
  • 209548 10/25/07 07:37 V. Sulkosky Loop 1 JT Opened 2 steps
  • 209547 10/25/07 07:10 V. Sulkosky Target --> Loop 1 He
  • 209545 10/25/07 06:53 Eric Voutier magnet screen
  • 209544 10/25/07 06:52 Eric Voutier Beam strip chart
  • 209543 10/25/07 06:31 V. Sulkosky VDC Effs. Run 20131
  • 209542 10/25/07 06:28 V. Sulkosky Large number of Zero Tracks
  • 209541 10/25/07 06:28 brads BPM + radiation monitor stripcharts
  • 209539 10/25/07 05:59 V. Sulkosky Target --> Beam Right Fe
  • 209536 10/25/07 05:24 V. Sulkosky Target --> Beam Right C
  • 209533 10/25/07 04:26 Eric Voutier Current limitation
  • 209531 10/25/07 04:17 V. Sulkosky Loop3 Beam Heat Load = 0
  • 209530 10/25/07 04:08 V. Sulkosky Target Stripcharts 2.6 uA on Loop3 LH + Pb
  • 209528 10/25/07 04:02 V. Sulkosky Target --> Loop3 LH + Pb
  • 209527 10/25/07 03:59 V. Sulkosky Right HV screens
  • 209526 10/25/07 03:56 V. Sulkosky Shower Checkout 3
  • 209524 10/25/07 03:55 V. Sulkosky Shower Checkout 2
  • 209523 10/25/07 03:53 V. Sulkosky Shower Checkout 1
  • 209522 10/25/07 03:52 V. Sulkosky Preshower Checkout 2
  • 209521 10/25/07 03:49 V. Sulkosky Preshower Checkout 1
  • 209520 10/25/07 03:43 V. Sulkosky Raster Size
  • 209519 10/25/07 03:37 V. Sulkosky Current limited to 6.8 uA
  • 209514 10/25/07 02:18 V. Sulkosky Target Screens with 6.8 uA on Loop1
  • 209512 10/25/07 02:09 V. Sulkosky Target --> He Loop 1
  • 209511 10/25/07 02:03 V. Sulkosky S2 (HRSR) Retiming Good
  • 209503 10/25/07 01:05 brads Mystery HV channels tripped in Beamline crate, first slot
  • 209498 10/25/07 00:45 jpchen Fan speed to 60 Hz, increased HP heater to be ready for 10 uA beam
  • 209497 10/25/07 00:11 jpchen Target partially vented during the day shift. Vacuum degraded.
  • 209495 10/25/07 00:09 Ramesh Shift Sumary
  • 209493 10/25/07 00:04 Chiranjib TARGET --->> OPTICS
  • Wednesday

  • 209491 10/24/07 23:48 Seonho Choi BCM rough calibration
  • 209490 10/24/07 23:38 Chiranjib TARGET --> Carbon ( right)
  • 209487 10/24/07 23:30 Chiranjib MORE ALARMs
  • 209486 10/24/07 23:25 Chiranjb Temp Sensor (loop 1)/ High power heater ( loop1)
  • 209485 10/24/07 23:22 Chiranjib UPPER LIMIT for loop 1 gas panel ( helium)
  • 209484 10/24/07 23:19 Chiranjib MORE ALARMS !!!!
  • 209483 10/24/07 23:13 Chiranjib TARGET ALARMS
  • 209482 10/24/07 23:06 brads Problems with OTR, BCMs, Harp 1H03B
  • 209481 10/24/07 23:05 brads Todo List for Wednesday SWING, Thursday OWL
  • 209480 10/24/07 22:00 Ramesh Getting pulsed beam
  • 209479 10/24/07 21:58 A. Camsonne HV Off on left arm
  • 209477 10/24/07 21:03 A. Camsonne Run 1228 long cosmics run
  • 209469 10/24/07 20:21 brads Masked out LHRS HVs in Alarm handler
  • 209468 10/24/07 20:17 brads Right P0 to 929 MeV/c
  • 209467 10/24/07 20:16 brads HV off for Left VDCs
  • 209466 10/24/07 20:13 Chiranjib TARGET SCREENS II
  • 209465 10/24/07 20:12 Chiranjib TARGET SCREENS I
  • 209464 10/24/07 18:22 Chiranjib SPOT ++
  • 209449 10/24/07 16:38 Chiranjib IOC reboot
  • 209447 10/24/07 16:01 Priyanga Target Screens
  • 209443 10/24/07 12:36 V. Sulkosky Added Delay to S1 (RHRS)
  • 209438 10/24/07 11:02 brads Ramped down Q1 on Left and Right arm
  • 209437 10/24/07 10:58 R. Michaels re: self-timing.
  • 209436 10/24/07 10:50 brads Left Q1 flow yellow (lo) alarm
  • 209435 10/24/07 10:42 brads Test from off-site
  • 209434 10/24/07 10:32 brads Unmasked LHRS HVs in the alarm handler
  • 209433 10/24/07 09:55 Priyanga Temperature jump
  • 209430 10/24/07 09:45 V. Sulkosky Timing Recheck with Cosmics (HRSR)
  • 209426 10/24/07 08:49 Priyanga Target Screens
  • 209423 10/24/07 08:05 Eric Voutier Shift summary
  • 209415 10/24/07 06:22 jpchen controlled access for NaI detector check out.
  • 209413 10/24/07 06:08 zhangyi Another run has been started
  • 209410 10/24/07 06:02 zhangyi Turn on the NaI detector and start run
  • 209407 10/24/07 05:10 zhangyi Run#20114 has been finished
  • 209405 10/24/07 04:38 zhangyi Target to Optics Target
  • 209404 10/24/07 04:01 jpchen Right arm detetcor check
  • 209402 10/24/07 03:25 jpchen moeved target to 12C earlier and moved to EMPTY now
  • 209400 10/24/07 03:16 jpchen spot++/spot_R not working
  • 209397 10/24/07 02:54 jpchen beam tuning completed, CW checkout , raster check now
  • Tuesday

  • 209394 10/23/07 23:04 Chiranjib TARGET SCREENS II
  • 209393 10/23/07 23:03 Chiranjib TARGET SCREENS I
  • 209392 10/23/07 23:02 V. Sulkosky Silver Calorimeter FSD Trip
  • 209391 10/23/07 23:02 Chiranjib ATARG alarm
  • 209389 10/23/07 20:34 brads Masked out LHRS HVs in the alarm handler
  • 209388 10/23/07 20:31 brads HV screens after LHRS HVs turned off
  • 209386 10/23/07 20:26 brads HV screens before turning off LHRS HVs
  • 209376 10/23/07 20:07 Xinhu Yan Database changed
  • 209375 10/23/07 19:51 brads Right Q1 doesn't respond to P0 setting
  • 209374 10/23/07 19:39 jpchen Cryotarget ready adn conditions
  • 209373 10/23/07 19:33 A. Camsonne Test weblogentry
  • 209372 10/23/07 19:19 f test
  • 209371 10/23/07 19:19 f f
  • 209370 10/23/07 19:13 brads Logbook web entry feature is broken...
  • 209369 10/23/07 19:11 brads If the printer won't work for you...
  • 209367 10/23/07 18:41 E.Chudakov HRSR shower/preshower status
  • 209366 10/23/07 18:38 brads Run plan: Tuesday SWING and Wednesday OWL
  • 209362 10/23/07 17:52 A. Camsonne Fan and HV on NaI detector for checks
  • 209358 10/23/07 17:15 V. Sulkosky S2 TDC Spectra (HRSR)
  • 209347 10/23/07 14:48 atrig Trigger-Download
  • 209344 10/23/07 14:36 atrig Trigger-Download
  • 209338 10/23/07 13:28 V. Sulkosky S2 TDC after Timing adjustments (RHRS)
  • 209317 10/23/07 10:25 atrig Trigger-Download
  • 209316 10/23/07 10:15 atrig Trigger-Download
  • 209315 10/23/07 10:11 atrig Trigger-Download
  • 209314 10/23/07 10:04 atrig Trigger-Download
  • 209313 10/23/07 09:41 atrig Trigger-Download
  • 209308 10/23/07 07:34 jpchen HVs and fans are off for NaI
  • 209307 10/23/07 07:33 jpchen NaI most channels have signals
  • 209292 10/23/07 06:21 jpchen NaI detector cabling completed
  • 209291 10/23/07 06:21 jpchen Hall not closed due to no RADCON checklist
  • 209289 10/23/07 05:45 A. Camsonne Trigger changed to OR of S1 S2 for NaI testing tonight
  • 209288 10/23/07 05:41 A. Camsonne HV 1440 on portserver hatsv4 port 6
  • Monday

  • 209287 10/22/07 20:20 brads Target ladder at EMPTY
  • 209286 10/22/07 18:51 Eugene Chuda HRSR preshower - a wide pedestal
  • Sunday

  • 209266 10/21/07 17:36 V. Sulkosky Light leaks Check S2 (HRSR)
  • Saturday

  • 209251 10/20/07 12:06 V. Sulkosky S2 and Cerenkov Centered (RHRS)
  • 209250 10/20/07 11:59 V. Sulkosky S1 ADC Channel Swap Fixed (RHRS)
  • 209247 10/20/07 09:22 V. Sulkosky Right Cerenkov HV Settings
  • 209246 10/20/07 09:20 G. Urciuoli/ HV adjustment Gas Cerenkov (RHSRS)
  • 209238 10/20/07 01:31 Yi Zhang/ ZE Detector channels to be checked
  • 209233 10/20/07 00:15 A. Camsonne FPP 1877 loaned to transversity
  • Friday

  • 209221 10/19/07 20:30 V. Sulkosky RHRS VDC Efficiencies and TDCs
  • 209220 10/19/07 20:25 V. Sulkosky HV Settings for S1 and S2 (RHRS)
  • 209219 10/19/07 20:23 V. Sulkosky S2 ADC Channels
  • 209218 10/19/07 20:16 V. Sulkosky HV adjustment for S2 (HRSR)
  • Thursday

  • 209154 10/18/07 22:51 V. Sulkosky HV adjustment for S1 (RHRS)
  • 209139 10/18/07 17:48 V. Sulkosky VDC after Connection Repairs
  • 209137 10/18/07 16:12 atrig Trigger-Download
  • Wednesday

    
    

    Tuesday

  • 209033 10/16/07 20:24 V. Sulkosky Right DAQ Scripts
  • Monday

  • 208966 10/15/07 16:42 V. Sulkosky Better Table format for S2 Cables
  • 208965 10/15/07 16:33 V. Sulkosky Recabling of S2 Part 2
  • Sunday

    
    

    Saturday

  • 208919 10/13/07 17:27 V. Sulkosky Recabling of S2
  • Friday

    
    

    Thursday

  • 208899 10/11/07 11:56 R. Michaels raster restoration
  • Tuesday

  • 208848 10/09/07 15:20 R. Michaels modified scaler.map
  • 208847 10/09/07 11:51 atrig Trigger-Download
  • 208846 10/09/07 11:31 atrig Trigger-Download
  • 208845 10/09/07 11:28 atrig Trigger-Download
  • 208844 10/09/07 11:23 atrig Trigger-Download
  • 208843 10/09/07 10:52 atrig Trigger-Download
  • Monday

  • 208838 10/08/07 16:13 atrig Trigger-Download
  • Saturday

  • 208836 10/06/07 13:52 atrig Trigger-Download
  • Friday

  • 208834 10/05/07 12:29 R. Michaels new overlap MLU
  • 208833 10/05/07 12:23 atrig Trigger-Download
  • 208832 10/05/07 12:22 atrig Trigger-Download
  • 208831 10/05/07 12:22 atrig Trigger-Download
  • 208830 10/05/07 12:17 atrig Trigger-Download
  • 208829 10/05/07 12:17 atrig Trigger-Download
  • Thursday

  • 208828 10/04/07 11:40 Higinbotham Test Entry
  • Wednesday

    
    

    Tuesday

  • 208826 10/02/07 14:50 atrig Trigger-Download